# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ Abstract.
We consider the stochastic dynamics of infinite particle systems in $\mathbb{R}^{d}$ interacting through the $d$-dimensional Coulomb potential. For arbitrary inverse temperature $\beta>0$ and all dimensions $d\geq 2$, we construct solutions to the associated infinite-dimensional stochastic differential equations (ISDEs). Our main results establish the existence of strong solutions and their pathwise uniqueness. The resulting labeled process is an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued diffusion with no invariant measure, while the associated unlabeled process is reversible with respect to the Coulomb random point field. Furthermore, we prove that finite-particle systems converge to these infinite-particle dynamics, thereby providing a rigorous foundation for Coulomb interacting Brownian motion.
Chubu University, osada@fsc.chubu.ac.jp
Kagoshima University, s-osada@edu.kagoshima-u.ac.jp
Keyword: Coulomb random point field infinite-dimensional stochastic differential equationinfinite-particle system
Acknowledgements: This work was supported by JSPS KAKENHI Grant Numbers JP21H04432, 22K13931
1. Introduction.
@@
$\Psi({x})$ $=\frac{1}{d-2}\frac{1}{|{x}|^{d-2}},\quad\text{ for }d\geq 3,$
$=-\log\lvert{x}\rvert,\quad\text{ for }d=2.$ (1.1)
@@
$\nabla\Psi({x})=-\frac{{x}}{|{x}|^{d}}.$ (1.2)
We call $\Psi$ the $d$-dimensional Coulomb potential. The object of this paper is to construct the stochastic dynamics of infinite-particle systems in $\mathbb{R}^{d}$ interacting through $\Psi$ at inverse temperature $\beta>0$. These dynamics are expected to satisfy the infinite-dimensional stochastic differential equation (ISDE)
$X_{t}^{i}-X_{0}^{i}=B_{t}^{i}+\frac{\beta}{2}\int_{0}^{t}\lim_{R\to\infty}\sum_{\lvert X_{u}^{i}-X_{u}^{j}\rvert<R,\ j\neq i}\frac{X_{u}^{i}-X_{u}^{j}}{\lvert X_{u}^{i}-X_{u}^{j}|^{d}}du\quad(i\in\mathbb{N}),$ (1.3)
@@
A striking feature of this problem is the exact correspondence between the ambient spatial dimension $d$ and the exponent in the Coulomb potential. This coincidence is not accidental: it is the source of both the depth and the difficulty of the problem. On the one hand, it ensures that Coulomb random point fields are natural equilibrium distributions in any dimension $d\geq 2$; on the other hand, it makes the analysis of the associated ISDE (1.3) extremely delicate, due to the long-range nature and singularity of the Coulomb force.
@@
To date, the only case where (1.3) has been solved is $d=\beta=2$ , corresponding to the Ginibre random point field. In this special situation, the determinantal structure of the equilibrium measure provides powerful tools, including sharp hyperuniform estimates for particle number fluctuations. Beyond $d=\beta=2$, however, such determinantal methods break down, and no general construction of the dynamics has been available.
@@
From the equilibrium viewpoint, the classical Gibbsian framework excludes Coulomb interactions because the potential is neither superstable nor integrable at infinity. The construction of Coulomb random point fields remained open for decades, until very recently it was resolved in general by Armstrong–Serfaty and Thoma [1, 44]. These works establish the existence of Coulomb random point fields for all $d\geq 2$ and $\beta>0$. The present paper addresses the complementary problem: the construction of the corresponding stochastic dynamics, i.e., reversible diffusions with respect to these measures.
Our approach is new, conceptually simple, and robust. For the first time, it provides a pathwise unique strong solution $\mathbf{X}$ of the Coulomb ISDE (1.3) for all spatial dimensions $d\geq 2$ and all inverse temperatures $\beta>0$. The central innovation of the method is the introduction of a new family of analytic estimates for the logarithmic derivatives of Coulomb random point fields. These estimates replace the determinantal and random matrix arguments that were previously available only in the Ginibre case. As a consequence, the solution $\mathbf{X}$ is established as an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued diffusion process without invariant measure, while the associated unlabeled process is shown to be a reversible diffusion with respect to the Coulomb random point field.
@@
We connect the finite ensembles to the ISDE through a two-step thermodynamic limit. For each radius $R$, we freeze the exterior configuration on $S_{R}^{c}=\{\lvert x\rvert\geq R\}$ and consider the $N$-particle gradient dynamics inside $S_{R}$ with normal reflection on $\partial S_{R}$. Denoting these finite-volume systems by $\mathbf{X}_{R}^{N}$, we first take the limit $N\to\infty$ at fixed $R$ to obtain an infinite-particle dynamics $\mathbf{X}_{R}$ in $S_{R}$ (reversible for the corresponding conditional measure). We then let $R\to\infty$ and prove $\mathbf{X}_{R}\Rightarrow\mathbf{X}$, where $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ is the unique strong solution to the Coulomb ISDE (1.3). The precise formulation of these statements, together with the assumptions required, is given in Section 2.
@@
In [28], the vanishing of self-diffusion was proved for the Ginibre interacting Brownian motion. The decisive mechanism behind this phenomenon is the number rigidity of the stationary random point field. With the present results in place, it is natural to formulate the following conjecture:
@@
We also note that Thoma [45] proved the absence of number rigidity for Coulomb random point fields for every $d\geq 3$, which strongly suggests that the self-diffusion coefficient is strictly positive in this regime. If confirmed, this dichotomy would vividly highlight the profound role of the long-range nature of the two-dimensional Coulomb potential, and firmly place Coulomb systems within the universality paradigm of singular interacting diffusions.
Organization of the paper.
The remainder of the paper is organized as follows. In Section 2, we state our main results and discuss related work and positioning. Section 3 presents examples. Section 45 establish the existence of the logarithmic derivative of the equilibrium measure and provide its explicit representation. Section 6 proves non-collision estimates for solutions of ISDEs. Section 79 introduce Dirichlet form techniques and construct weak solutions under both lower and upper schemes. Section 10 establishes the existence of unique strong solutions, while Section 11 demonstrates the uniqueness of Dirichlet forms and completes the proof of our diffusion theorem. Section 12 shows the convergence of finite-particle dynamics to the infinite-particle ISDE dynamics. Finally, the Appendix (Section 13) collects auxiliary results on configuration spaces, Dirichlet forms, and the general theory of ISDEs.
2. Main theorems (Theorem 2.1–Theorem 2.5)
We now formulate the problem in detail and state our main results.
@@ 2.1. Set up and logarithmic derivatives
$\mathsf{S}=\Bigl{\{}\mathsf{s}\,;\,\mathsf{s}=\sum_{i}\delta_{s^{i}},\mathsf{s}(S_{R})<\infty\quad\text{ for all }R\in\mathbb{N}\Bigr{\}}.$ (2.1)
@@
For $R\in\mathbb{N}$ and $m\in\{0\}\cup\mathbb{N}$, set $\mathsf{S}_{R}^{m}=\{\mathsf{s}\in\mathsf{S};\mathsf{s}(S_{R})=m\}$. Then, for each $R\in\mathbb{N}$,
$\mathsf{S}=\bigsqcup_{m=0}^{\infty}\mathsf{S}_{R}^{m}.$ (2.2)
For $A\in\mathscr{B}(\mathsf{S})$, define $\pi_{A}(\mathsf{s})=\mathsf{s}(\cdot\cap A)$. Let $S_{R}^{m}$ denote the $m$-fold product of $S_{R}$. For $\mathsf{s}\in\mathsf{S}_{R}^{m}$, an $S_{R}^{m}$-coordinate of $\mathsf{s}$ is a tuple $\mathbf{x}_{R}^{m}(\mathsf{s})\in S_{R}^{m}$ such that
$\pi_{S_{R}}(\mathsf{s})=\sum_{i=1}^{m}\delta_{x_{R}^{i}(\mathsf{s})},\qquad\mathbf{x}_{R}^{m}(\mathsf{s})=(x_{R}^{i}(\mathsf{s}))_{i=1}^{m}.$
Definition 2.1.
For $f:\mathsf{S}\to\mathbb{R}$ and $R,m\in\mathbb{N}$, let $f_{R,\mathsf{s}}^{m}(\mathbf{x})$ be the function defined by
$f_{R,\cdot}^{m}:\mathsf{S}\times S_{R}^{m}\to\mathbb{R},\qquad(\mathsf{s},\mathbf{x})\mapsto f_{R,\mathsf{s}}^{m}(\mathbf{x}),$ (2.3)
subject to the following conditions:
$f_{R,\mathsf{s}}^{m}(\mathbf{x})\ \text{is symmetric in $\mathbf{x}$ for each $\mathsf{s}\in\mathsf{S}_{R}^{m}$},$ (2.4)
$f_{R,\mathsf{s}(1)}^{m}(\mathbf{x})=f_{R,\mathsf{s}(2)}^{m}(\mathbf{x})\ \text{if }\pi_{S_{R}^{c}}(\mathsf{s}(1))=\pi_{S_{R}^{c}}(\mathsf{s}(2)),$ (2.5)
$f_{R,\mathsf{s}}^{m}(\mathbf{x}_{R}^{m}(\mathsf{s}))=f(\mathsf{s})\ \text{for }\mathsf{s}\in\mathsf{S}_{R}^{m},$ (2.6)
$f_{R,\mathsf{s}}^{m}(\mathbf{x})=0\ \text{for }\mathsf{s}\notin\mathsf{S}_{R}^{m}.$ (2.7)
$f(\mathsf{s})=\sum_{m=0}^{\infty}f_{R,\mathsf{s}}^{m}(\mathbf{x}_{R}^{m}(\mathsf{s})).$
By convention, $\mathbf{x}_{R}^{0}(\mathsf{s})=\emptyset$ for $\mathsf{s}\in\mathsf{S}_{R}^{0}$, and $f_{R,\mathsf{s}}^{0}$ is constant on $\mathsf{S}_{R}^{0}$. For a bounded set $A$, one defines $\mathbf{x}_{A}^{m}(\mathsf{s})$ and $f_{A,\mathsf{s}}^{m}(\mathbf{x})$ analogously by replacing $S_{R}$ with $A$.
A function $f:\mathsf{S}\to\mathbb{R}$ is called smooth if $f_{R,\mathsf{s}}^{m}(\mathbf{x})$ is smooth in $\mathbf{x}$ on $S_{R}^{m}$ for all $R,m,\mathsf{s}$, and is called local if $f$ is $\sigma[\pi_{S_{R}}]$-measurable for some $R\in\mathbb{N}$. We write
@@
$\mathscr{D}_{\bullet\mathrm{b}}$ $=\{f\in\mathscr{D}_{\bullet}\,;\,f\ \text{is bounded}\},\qquad\mathscr{D}_{\circ\mathrm{b}}=\{f\in\mathscr{D}_{\circ}\,;\,f\ \text{is bounded}\}.$ (2.8)
Remark 2.1.
Even if $f$ is smooth, it need not be $\mathscr{B}(\mathsf{S})$-measurable. The smoothness above refers to partial smoothness in $\mathbf{x}$ for each fixed $\mathsf{s}$.
A probability measure $\nu$ on $(\mathsf{S},\mathscr{B}(\mathsf{S}))$ is called a random point field on $\mathbb{R}^{d}$. A symmetric, locally integrable function $\rho_{\nu}^{m}$ on $(\mathbb{R}^{d})^{m}$ is called the $m$-point correlation function of $\nu$ (with respect to Lebesgue measure) if
$\int_{A_{1}^{k_{1}}\times\cdots\times A_{n}^{k_{n}}}\rho_{\nu}^{m}({x}^{1},\ldots,{x}^{m})\,d{x}^{1}\cdots d{x}^{m}=\int_{\mathsf{S}}\prod_{i=1}^{n}\frac{\mathsf{s}(A_{i})!}{(\mathsf{s}(A_{i})-k_{i})!}\,\nu(d\mathsf{s}),$ (2.9)
for any disjoint bounded measurable sets $A_{1},\ldots,A_{n}$ and integers $k_{1}+\cdots+k_{n}=m$. When $\mathsf{s}(A_{i})-k_{i}<0$, we interpret the factorial as zero. Here, $\mathsf{s}(A)$ denotes the number of particles in $A$ when $\mathsf{s}=\sum_{i}\delta_{s^{i}}$.
For a random point field $\nu$ enjoying the one-point correlation function $\rho_{\nu}^{1}$, we define the one-reduced Campbell measure $\nu^{[1]}$ of $\nu$ by
$\nu^{[1]}(dxd\mathsf{s})=\rho_{\nu}^{1}(x)\nu_{x}(d\mathsf{s})dx,$
where $\nu_{x}=\nu(\cdot+\delta_{x}|\mathsf{s}(\{{x}\})\geq 1)$ is the reduced Palm measure of $\nu$ conditioned at ${x}\in\mathbb{R}^{d}$. We set $L_{\mathrm{loc}}^{1}(\nu^{[1]})=\bigcap_{R=1}^{\infty}L^{1}(\nu^{[1]}(\cdot\cap S_{R}\times\mathsf{S}))$. The concept of one-reduced Campbell measure can be generalized to $m$-reduced Campbell measure and for $\nu$ not necessarily enjoying factorial moments (see [12]).

### v4

# Infinite-dimensional stochastic differential equations for Coulomb random point fields

@@ Abstract.
We study the infinite-dimensional stochastic differential equations (ISDEs) of infinite-particle systems associated with Coulomb random point fields. The stochastic dynamics described by these ISDEs are referred to as Coulomb interacting Brownian motions. In all spatial dimensions $d\geq 2$ and for all inverse temperatures $\beta>0$, we construct the Coulomb interacting Brownian motions.
We prove that the ISDEs admit strong solutions and that pathwise uniqueness holds. The resulting labeled dynamics form an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued diffusion, possibly without an invariant measure, while the corresponding unlabeled process is a reversible diffusion with respect to the underlying Coulomb random point field.
Moreover, we identify the infinite-particle stochastic dynamics as the limit in path space of finite-particle systems driven by stochastic differential equations. This identification is achieved through two approximation schemes: finite-domain systems with reflecting boundary conditions and $N$-particle systems. Although the $N$-particle approximation is more fundamental, its justification relies crucially on the finite-domain approximation together with the uniqueness of solutions to the ISDEs.
Previously, only the case $d=2$ and $\beta=2$, known as the Ginibre interacting Brownian motion, was understood through random matrix theory and determinantal random point fields. Extending this result beyond the determinantal setting has remained a major difficulty.
We introduce a new, conceptually clear method based on stochastic analysis of infinite-particle systems with long-range interactions that yields a rigorous construction of Coulomb interacting Brownian motions. A key ingredient is an explicit computation of the logarithmic derivatives of Coulomb random point fields.
keywords: Coulomb random point fields; stochastic dynamics of infinite-particle systems; logarithmic derivatives of random point fields; infinite-dimensional stochastic differential equations
Chubu University, osada "AT"fsc.chubu.ac.jp
Kagoshima University, s-osada"AT"edu.kagoshima-u.ac.jp
H. Osada, Chubu University, osada@fsc.chubu.ac.jp
S. Osada, Kagoshima University, s-osada@edu.kagoshima-u.ac.jp
1. Introduction
The aim of this paper is to develop a theory of infinite-dimensional stochastic differential equations (ISDEs) arising from long-range interacting particle systems. We focus on the construction and characterization of strong solutions, their pathwise uniqueness, and the relationship between labeled and unlabeled dynamics within the framework of stochastic analysis on infinite-particle systems. Moreover, we derive the limiting stochastic dynamics as the limit of $N$-particle systems described by solutions of finite-dimensional SDEs.
@@ 1. Introduction
$\Psi({x})=\begin{cases}\dfrac{1}{d-2}\dfrac{1}{|{x}|^{d-2}}&d\geq 3,\\[6.0pt] -\log\lvert{x}\rvert&d=2.\end{cases}$ (1.1)
@@
$\nabla\Psi({x})=-\frac{{x}}{|{x}|^{d}}.$
We call $\Psi$ the $d$-dimensional Coulomb potential. The purpose of this paper is to construct the stochastic dynamics of infinite-particle systems in $\mathbb{R}^{d}$ interacting through $\Psi$ at inverse temperature $\beta>0$.
These dynamics are expected to satisfy an ISDE, whose prototypical form is
$X_{t}^{i}-X_{0}^{i}=B_{t}^{i}+\frac{\beta}{2}\int_{0}^{t}\lim_{R\to\infty}\sum_{\lvert X_{u}^{i}-X_{u}^{j}\rvert<R,\ j\neq i}\frac{X_{u}^{i}-X_{u}^{j}}{\lvert X_{u}^{i}-X_{u}^{j}|^{d}}du,\quad i\in\mathbb{N},$ (1.2)
@@
Equilibrium states of the unlabeled dynamics of Coulomb interacting Brownian motion are given by Coulomb random point fields (RPFs).
Coulomb RPFs are probability measures describing unlabeled particle systems with interaction potential $\Psi$. They are obtained as limits of finite-particle systems whose labeled distributions $\widetilde{\mu}^{N}$ are given by the density
$\mathfrak{m}^{N}(\mathbf{x})=\frac{1}{\mathscr{Z}}\exp\Big\{-\beta\Big(\sum_{i=1}^{N}\Phi^{N}({x}^{i})+\sum_{i<j}^{N}\Psi^{N}({x}^{i},{x}^{j})\Big)\Big\},$ (1.3)
where $0<\beta<\infty$, $\Phi^{N}$ denotes a confining potential, and $\Psi^{N}$ an interaction potential at the $N$-particle level, which may coincide with $\Psi$.
We assume that these potentials converge to limiting potentials:
$\lim_{N\to\infty}\nabla\Phi^{N}({x})=\nabla\Phi({x}),\qquad\lim_{N\to\infty}\Psi^{N}({x},{y})=\Psi({x}-{y}).$ (1.4)
We refer to ((A1)) for the precise meaning of (1.4).
The $N$-particle dynamics $\mathbf{X}^{N}=(X^{N,i})_{i=1}^{N}$ are gradient dynamics associated with the Dirichlet form $\widetilde{\mathscr{E}}^{N}$ on $L^{2}(\widetilde{\mu}^{N})$, which is defined by
$\widetilde{\mathscr{E}}^{N}(f,g)=\int_{(\mathbb{R}^{d})^{N}}\frac{1}{2}\sum_{i=1}^{N}\big(\nabla_{{x}^{i}}f,\nabla_{{x}^{i}}g\big)_{\mathbb{R}^{d}}\,\widetilde{\mu}^{N}(d\mathbf{x}).$
Here we denote by $(\cdot,\cdot)_{\mathbb{R}^{d}}$ the Euclidean inner product on $\mathbb{R}^{d}$.
By integration by parts, we have
$\widetilde{\mathscr{E}}^{N}(f,g)$ $=-\int_{(\mathbb{R}^{d})^{N}}\frac{1}{2}\sum_{i=1}^{N}\Big\{\Delta_{{x}^{i}}f+\big((\nabla_{{x}^{i}}\log\mathfrak{m}^{N}),\nabla_{{x}^{i}}f\big)_{\mathbb{R}^{d}}\Big\}g\,\widetilde{\mu}^{N}(d\mathbf{x}).$
See Subsection 13.2 for the definition of Dirichlet forms used here. By (1.3),
$\nabla_{{x}^{i}}\log\mathfrak{m}^{N}(\bm{x})=-\beta\Big(\nabla\Phi^{N}({x}^{i})+\sum_{\begin{subarray}{c}j\neq i\\ 1\leq j\leq N\end{subarray}}\nabla_{{x}}\Psi^{N}({x}^{i},{x}^{j})\Big).$
Consequently, $\mathbf{X}^{N}$ satisfies the following SDE in differential form:
$dX_{t}^{N,i}$ $=dB_{t}^{i}+\frac{\beta}{2}\Big(-\nabla\Phi^{N}(X_{t}^{N,i})-\sum_{\begin{subarray}{c}j\neq i\\ 1\leq j\leq N\end{subarray}}\nabla_{{x}}\Psi^{N}(X_{t}^{N,i},X_{t}^{N,j})\Big)\,dt.$ (1.5)
Taking the limit $N\to\infty$, we formally obtain the limiting ISDE:
$dX_{t}^{i}$ $=dB_{t}^{i}+\frac{\beta}{2}\Big(-\nabla\Phi(X_{t}^{i})+\lim_{R\to\infty}\sum_{\begin{subarray}{c}j\neq i\\ \lvert X_{t}^{j}\rvert<R\end{subarray}}\frac{X_{t}^{i}-X_{t}^{j}}{\lvert X_{t}^{i}-X_{t}^{j}\rvert^{d}}\Big)\,dt,\quad i\in\mathbb{N}.$ (1.6)
If, in addition, $\mu$ is translation invariant, this equation reduces to the ISDE (1.2). Such multiple representations of the stochastic dynamics arise from the long-range nature of the Coulomb interaction.
@@
A striking feature of this problem is the exact correspondence between the ambient spatial dimension $d$ and the exponent in the Coulomb potential. This coincidence is not accidental: it is the source of both the depth and the difficulty of the problem. On the one hand, it ensures that Coulomb random point fields (RPFs) are natural equilibrium distributions in any dimension $d\geq 2$; on the other hand, it makes the analysis of the associated ISDE (1.6) extremely delicate, due to the long-range nature and singularity of the Coulomb force.
@@
To date, the only case where (1.6) has been solved is $d=\beta=2$, corresponding to the Ginibre RPF. In this special situation, the determinantal structure of the equilibrium measure provides powerful tools, including sharp hyperuniform estimates for particle number fluctuations. Beyond $d=\beta=2$, however, such determinantal methods break down, and no general construction of the dynamics has been available.
@@
From the equilibrium viewpoint, the classical Gibbsian framework excludes Coulomb interactions because the potential is neither superstable nor integrable at infinity. The construction of Coulomb RPFs remained open for decades, until very recently it was resolved in general by Armstrong–Serfaty and Thoma [1, 44]. These works establish the existence of Coulomb RPFs for all $d\geq 2$ and $\beta>0$. The present paper addresses the complementary problem: the construction of the corresponding stochastic dynamics, i.e., reversible diffusions with respect to these measures.
Our approach is new, conceptually simple, and robust.
For the first time, it constructs a stochastic dynamics for infinite-particle Coulomb systems in all spatial dimensions $d\geq 2$ and for all inverse temperatures $\beta>0$ (Theorem 2.2).
Moreover, this dynamics is realized as a pathwise unique strong solution of the Coulomb ISDE (1.6) with explicit coefficients (Theorem 2.3). This follows from the explicit representation of the logarithmic derivative of $\mu$ (Theorem 2.1).
The central innovation of the method is the introduction of a new family of analytic estimates for the logarithmic derivatives of Coulomb RPFs. These estimates replace the determinantal and random matrix arguments that were previously available only in the Ginibre case.
The solution $\mathbf{X}$ is constructed as an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued diffusion process without relying on an invariant or reference measure, while the associated unlabeled process is shown to be a reversible diffusion with respect to the Coulomb RPF (Theorem 2.4).
We derive the ISDE under two different thermodynamic limits.
One approach is based on finite-domain approximations, in which finite-particle systems are connected through a two-step thermodynamic limit (Theorem 12.1).
@@
For each radius $R$, we freeze the exterior configuration on $\overline{S}_{R}^{c}=\{\lvert s\rvert>R\}$ and consider the $N$-particle gradient dynamics inside $S_{R}=\{\lvert s\rvert<R\}$ with normal reflection on $\partial S_{R}$. Denoting these finite-volume systems by $\mathbf{X}_{R}^{N}$, we first take the limit $N\to\infty$ with $R$ fixed to obtain an infinite-particle dynamics $\mathbf{X}_{R}$ in $S_{R}$, which is reversible with respect to the corresponding conditional measure. We then let $R\to\infty$ and prove that $\mathbf{X}_{R}$ converges in law to $\mathbf{X}$, where $\mathbf{X}$ is the unique strong solution to the Coulomb ISDE (1.6).
The other approach is based on $N$-particle approximations. In this approach, solutions $\mathbf{X}^{N}$ of the finite-dimensional SDE (1.5) converge in the path space to the solution $\mathbf{X}$ of the ISDE (Theorem 2.5).
More precisely, we regard $\mathbf{X}^{N}$ as an $(\mathbb{R}^{d})^{\mathbb{N}}$-valued process by introducing frozen dummy particles for each $i>N$, which do not affect the dynamics.
In Theorem 2.5, we establish that
$\mathbf{X}=\lim_{N\to\infty}\mathbf{X}^{N}\quad\text{in law in }C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}.$
Although the second derivation is the most fundamental, it relies on the first derivation, on the convergence of a related family of Dirichlet forms, and on a sandwich-type convergence theorem built upon these results. The uniqueness of solutions to the ISDE is indispensable for implementing the sandwich-type convergence.
These results justify the ISDE and provide a simulation scheme for the limiting stochastic dynamics.
@@
In [31], the vanishing of self-diffusion was proved for the Ginibre interacting Brownian motion. The decisive mechanism behind this phenomenon is the number rigidity of the stationary RPF. In contrast, in [25], it was proved that the self-diffusion coefficient is strictly positive for any potential in Ruelle’s class with convex hardcore in dimensions $d\geq 2$.
Taken together, these results highlight a structural distinction between Coulomb interactions and Ruelle-type interactions. With the present results in place, it is natural to formulate the following conjecture:
@@
In [45], Thoma proved number rigidity for Coulomb RPFs in the case $d=2$, and established the absence of number rigidity for Coulomb RPFs in all dimensions $d\geq 3$. This dichotomy strongly suggests that the self-diffusion coefficient is strictly positive in dimensions $d\geq 3$.
If confirmed, this phenomenon would vividly highlight the profound role played by the long-range nature of the two-dimensional Coulomb potential, and would firmly place Coulomb systems within the universality paradigm of singular interacting diffusions.
Organization of the paper. The remainder of the paper is organized as follows. In Section 2, we state our main results and discuss related work and positioning. Section 3 presents examples. Sections 4 and 5 establish the existence of the logarithmic derivative of the equilibrium measure and provide its explicit representation. Section 6 proves non-collision estimates for solutions of ISDEs. Sections 79 introduce Dirichlet form techniques and construct weak solutions under both lower and upper schemes. Section 10 establishes the existence of unique strong solutions, while Section 11 demonstrates the uniqueness of Dirichlet forms and completes the proof of our diffusion theorem. Section 12 shows the convergence of finite-particle dynamics to the infinite-particle ISDE dynamics. Finally, the Appendix (Section 13) collects auxiliary results on configuration spaces, Dirichlet forms, and the general theory of ISDEs.
2. Setup and main theorems
We now formulate the problem in detail and state our main results (Theorems 2.12.5).
Throughout the paper, we occasionally attach labels directly to constants to facilitate later references. These labels are independent of equation numbering and are used consistently whenever such constants reappear.
Moreover, when an unlabeled configuration $\mathsf{s}$ is written together with points $s^{i}$, we implicitly use a labeled representation $\mathsf{s}=\sum_{i}\delta_{s^{i}}$.
@@ 2.1. Set up and logarithmic derivatives
$\mathsf{S}=\Bigl\{\mathsf{s}\,;\,\mathsf{s}=\sum_{i}\delta_{s^{i}},\mathsf{s}(S_{R})<\infty\quad\text{ for all }R\in\mathbb{N}\Bigr\}.$
@@
A function $f:\mathsf{S}\to\mathbb{R}$ is called smooth if $f_{R,\mathsf{s}}^{m}$ is smooth on $S_{R}^{m}$ for all $R,m\in\mathbb{N}$ and $\mathsf{s}$, where $f_{R,\mathsf{s}}^{m}$ is the $S_{R}^{m}$-representation of $f$ defined in Definition 13.6. It is called local if $f$ is $\sigma[\pi_{S_{R}}]$-measurable for some $R\in\mathbb{N}$, where $\pi_{A}(\mathsf{s})=\mathsf{s}(\cdot\cap A)$ for $A\subset\mathbb{R}^{d}$. We write
@@
$\mathscr{D}_{\bullet\mathrm{b}}$ $=\{f\in\mathscr{D}_{\bullet}\,;\,f\ \text{is bounded}\},\quad\mathscr{D}_{\circ\mathrm{b}}=\{f\in\mathscr{D}_{\circ}\,;\,f\ \text{is bounded}\}.$
A probability measure $\nu$ on $(\mathsf{S},\mathscr{B}(\mathsf{S}))$ is called an RPF on $\mathbb{R}^{d}$. For an RPF $\nu$ with the one-point correlation function $\rho_{\nu}^{1}$, we define the one-reduced Campbell measure $\nu^{[1]}$ of $\nu$ by
$\nu^{[1]}(dxd\mathsf{s})=\rho_{\nu}^{1}(x)\nu_{x}(d\mathsf{s})dx,$ (2.1)
where $\nu_{x}=\nu(\cdot+\delta_{x}|\mathsf{s}(\{{x}\})\geq 1)$ is the reduced Palm measure of $\nu$ conditioned at ${x}\in\mathbb{R}^{d}$. We set $L_{\mathrm{loc}}^{1}(\nu^{[1]})=\bigcap_{R=1}^{\infty}L^{1}(\nu^{[1]}(\cdot\cap S_{R}\times\mathsf{S}))$. We refer to Definition 13.5 for the definition of correlation functions.
Let $\mathscr{D}_{\circ\mathrm{b}}^{[1]}:=C_{0}^{\infty}(\mathbb{R}^{d})\otimes\mathscr{D}_{\circ\mathrm{b}}$ and $\mathscr{D}_{\bullet\mathrm{b}}^{[1]}:=C_{0}^{\infty}(\mathbb{R}^{d})\otimes\mathscr{D}_{\bullet\mathrm{b}}$.